OpenStudy (anonymous):

Write an equation of the line given the following conditions: The line that contains the endpoints (5,-3) & (11,9).

OpenStudy (anonymous):

Ok, do you have any idea how to start the problem?

OpenStudy (anonymous):

to be honest no

OpenStudy (anonymous):

I can make a draw. 

OpenStudy (anonymous):

Do you need to use an specific method?

OpenStudy (anonymous):

in y=mx+b form

OpenStudy (anonymous):

You know that those points satisfy the equation of the line. So you can plug them in to that equation.

OpenStudy (anonymous):

for the first point: -3 = m*5 + b

OpenStudy (anonymous):

Can you do the same for the second point?

OpenStudy (anonymous):

I found the slope I just need to figure out the y-intercept

OpenStudy (anonymous):

Ohh great, what did you get for the slope?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

wel you can substitute that number into the equation I wrote: -3 = (2)*5 + b

OpenStudy (anonymous):

and solve for b

OpenStudy (anonymous):

y=2x-13 is that the correct equation

OpenStudy (anonymous):

You can check by plugging the coordinates of the points into the equation.

OpenStudy (anonymous):

It seems it works for the first point.

OpenStudy (anonymous):

thanks it worked but I have another 2 questions can you help me with them

OpenStudy (anonymous):

And it works for the second, great!

OpenStudy (anonymous):

Okay, lets go!

OpenStudy (anonymous):

Did you post it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't see it.